OpenStudy (anonymous):

Consider the line . What is the slope of a line parallel to this line? What is the slope of a line perpendicular to this line?

OpenStudy (aroub):

whats the line?

OpenStudy (anonymous):

x-3y=7

OpenStudy (aroub):

parallel slope=1/3

OpenStudy (aroub):

perpendicular slope= -3

OpenStudy (anonymous):

thank you so much how do you do it ( a basic summary is fine)

OpenStudy (aroub):

okay..

OpenStudy (aroub):

u return it to form y=mx+b where m is the slope, y=1/3x-7/3 parallel slope=same slope so on here u have 1/3 perpendicular slope= opposite reciprocal of the slope u have the reciprocal of 1/3 is 3 and the opposite of 3 is -3